### v1

# Numerical solution of elliptic distributed optimal control problems with boundary value tracking

Abstract
We consider some boundary value tracking optimal control problem constrained by a Neumann boundary value problem for some elliptic partial differential equation where the control acts as right-hand side. This optimal control problem can be reformulated as a state-based variational problem that is the starting point for the finite element discretizion. In this paper, we only consider a tensor-product finite element discretizion for which optimal discretization error estimates and fast solvers can be derived. Numerical experiments illustrate the theoretical results quantitatively.
1 Introduction
Let us consider the following boundary tracking optimal control problem (ocp): Find the state $y_{\varrho}\in H^{1}(\Omega)$ and the optimal control $u_{\varrho}\in U$ minimizing the cost functional
$J(y_{\varrho},u_{\varrho})=\frac{1}{2}\,\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}^{2}+\frac{1}{2}\,\varrho\,\|u_{\varrho}\|_{U}^{2}$ (1)
subject to the Neumann boundary value problem (bvp)
$\Delta y_{\varrho}+y_{\varrho}=u_{\varrho}\quad\mbox{in}\;\Omega,\quad\partial_{n}y_{\varrho}=0\quad\mbox{on}\;\Gamma,$ (2)
where $\overline{y}\in L^{2}(\Gamma)$ denotes a given target, $\varrho\in{\mathbb{R}}_{+}$ is a positive regularization or cost parameter, and $\Omega\subset\mathbb{R}^{d}$, $d=2,3$, is a bounded Lipschitz domain with the boundary $\Gamma=\partial\Omega$. Our work was inspired by the paper [4], where the $L^{2}$ regularization corresponding to the choice $U=L^{2}(\Omega)$ was investigated. The boundary tracking is a special case of the partial tracking of a given target in a subset of $\Omega$ that is also called limited observation; see, e.g., [5]. In this contribution, we consider the energy regularization corresponding to the choice $U=\widetilde{H}^{-1}(\Omega):=[H^{1}(\Omega)]^{*}$, where we use the standard notations for Lebesgue and Sobolev spaces; see, e.g., [8]. In order to follow the abstract theory presented in [3], we define the state space in such a way that the state-to-control map is an isomorphism. This allows us to derive a state-based formulation which is the basis for the numerical solution. Here we restrict the analysis to a conforming tensor-product finite element (fe) discretization that finally leads to a linear system of algebraic equations for which fast solvers can be constructed.
2 State-based variational reformulation
The variational formulation of the Neumann bvp (2) reads to find $y_{\varrho}\in H^{1}(\Omega)$ such that
$\langle\nabla y_{\varrho},\nabla y\rangle_{L^{2}(\Omega)}+\langle y_{\varrho},y\rangle_{L^{2}(\Omega)}=\langle u_{\varrho},y\rangle_{\Omega}$ (3)
is satisfied for all $y\in H^{1}(\Omega)$, where we assume $u_{\varrho}\in\widetilde{H}^{-1}(\Omega)$. While the Neumann boundary condition in (2) enters the variational formulation (3) in a natural way, this condition has to be included in the definition of the state space
$Y:=\Big\{y\in H^{1}(\Omega):\langle\partial_{n}y,\phi\rangle_{\Gamma}=0\;\mbox{for all}\;\phi\in H^{1/2}(\Gamma)\Big\}.$
When using duality arguments, we then conclude
$\|u_{\varrho}\|_{Y^{*}}:=\sup\limits_{0\neq y\in Y}\frac{\langle u_{\varrho},y\rangle_{\Omega}}{\|y\|_{H^{1}(\Omega)}}=\sup\limits_{0\neq y\in Y}\frac{\langle y_{\varrho},y\rangle_{H^{1}(\Omega)}}{\|y\|_{H^{1}(\Omega)}}=\|y_{\varrho}\|_{H^{1}(\Omega)},$
and instead of (1) we can consider the reduced state-based cost functional
$\widetilde{J}(y_{\varrho})=\frac{1}{2}\,\|y_{\varrho}-\overline{y}\|^{2}_{L^{2}(\Gamma)}+\frac{1}{2}\,\varrho\,\|y_{\varrho}\|^{2}_{H^{1}(\Omega)},$ (4)
whose minimizer $y_{\varrho}\in Y$ is the unique solution of the gradient equation satisfying
$\langle y_{\varrho},y\rangle_{L^{2}(\Gamma)}+\varrho\,\langle y_{\varrho},y\rangle_{H^{1}(\Omega)}=\langle\overline{y},y\rangle_{L^{2}(\Gamma)}\quad\mbox{for all}\;y\in Y.$ (5)
Following the abstract theory as given in [3, Lemma 2.1], we have the following regularization error estimate.
Lemma 2.1.
Let $y_{\varrho}\in Y$ be the unique solution of the variational formulation (5). For $\overline{y}\in L^{2}(\Gamma)$ there hold the estimates
$\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}\leq\|\overline{y}\|_{L^{2}(\Gamma)},\quad\|y_{\varrho}\|_{H^{1}(\Omega)}\leq\varrho^{-1/2}\,\|\overline{y}\|_{L^{2}(\Gamma)}.$ (6)
If  $\overline{y}\in H^{1/2}(\Gamma)$ is the Dirichlet trace of  $\overline{y}_{e}\in Y$, then $\|y_{\varrho}\|_{H^{1}(\Omega)}\leq\|\overline{y}_{e}\|_{H^{1}(\Omega)}$, and
$\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}\leq\varrho^{1/2}\,\|\overline{y}_{e}\|_{H^{1}(\Omega)},\quad\|y_{\varrho}-\overline{y}_{e}\|_{H^{1}(\Omega)}\leq\|\overline{y}_{e}\|_{H^{1}(\Omega)}.$ (7)
If $\overline{y}\in H^{1}(\Gamma)$ is the Dirichlet trace of $\overline{y}_{e}\in Y\cap H^{3/2+\varepsilon}(\Omega)$ for some $\varepsilon>0$, then
$\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}\leq c\,\varrho\,\|\overline{y}\|_{H^{1}(\Gamma)},\quad\|y_{\varrho}-\overline{y}_{e}\|_{H^{1}(\Omega)}\leq c\,\varrho^{1/2}\,\|\overline{y}\|_{H^{1}(\Gamma)}.$ (8)
Note that the extension $\overline{y}_{e}\in Y$ of $\overline{y}\in H^{1/2}(\Gamma)$ as used in Lemma 2.1 can be computed as solution of the first boundary value problem for the biharmonic equation,
$\Delta^{2}\overline{y}_{e}=0\quad\mbox{in}\;\Omega,\quad\overline{y}_{e}=\overline{y},\quad\partial_{n}\overline{y}_{e}=0\quad\mbox{on}\;\Gamma.$ (9)
3 Conforming FE discretization on tensor product meshes
For a conforming fe discretization of the variational formulation (5), we need to introduce a fe space $Y_{h}\subset Y$ of basis functions with zero normal derivatives. In this paper, we restrict our considerations to the unit square $\Omega=(0,1)^{d}$ which allows us to use appropriate tensor product fe spaces. Let $\widetilde{S}_{h}^{1}(0,1):=\mbox{span}\{\varphi_{i}\}_{i=1}^{n-1}$ be the modified fe space of piecewise linear and continuous basis functions $\varphi_{i}$ which are defined with respect to a decomposition $0=x_{0}<x_{1}<x_{2}<\ldots<x_{n-1}<x_{n}=1$ of the unit interval, with the local mesh sizes $h_{i}:=x_{i}-x_{i-1}$, $i=1,\ldots,n$, and with the global mesh size $h:=\max_{i}h_{i}$. While the basis functions $\varphi_{i}$ for $i=2,\ldots,n-2$ are the standard piecewise linear and continuous ones, the basis functions $\varphi_{1}$ and $\varphi_{n-1}$ are $1$ in the intervals $(x_{0},x_{1})$ and $(x_{n-1},x_{n})$, respectively; see Fig. 1.
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$x_{n-3}$$x_{n-2}$$x_{n-1}$$x_{n}$$\varphi_{2}$$\varphi_{n-2}$$\varphi_{1}$$\varphi_{n-1}$
Figure 1: Modified piecewise linear basis functions $\varphi_{i}(x)$, $i=1,\ldots,n-1$.
By construction, we have $\varphi_{i}^{\prime}(x)=0$ for $x\in\{0,1\}$, $i=1,\ldots,n-1$. We now define the conforming fe space $Y_{h}=\otimes_{i=1}^{d}\widetilde{S}_{h}^{1}(0,1)=\mbox{span}\{\phi_{k}\}_{k=1}^{m}\subset Y$ of piecewise multi-linear continuous basis functions $\phi_{k}$ with vanishing Neumann trace $\partial_{n}\phi_{k}$ on $\Gamma$. We note that $m=(n-1)^{d}$. The fe discretization of the variational formulation (5) leads to the finite element scheme: Find $y_{\varrho h}\in Y_{h}$ such that
$\langle y_{\varrho h},y_{h}\rangle_{L^{2}(\Gamma)}+\varrho\,\langle y_{\varrho h},y_{h}\rangle_{H^{1}(\Omega)}=\langle\overline{y},y_{h}\rangle_{L^{2}(\Gamma)}\quad\mbox{for all}\;y_{h}\in Y_{h}.$ (10)
Using standard arguments, we immediately arrive at the Cea-type estimate
$\|y_{\varrho}-y_{\varrho h}\|_{L^{2}(\Gamma)}^{2}+\varrho\,\|y_{\varrho}-y_{\varrho h}\|^{2}_{H^{1}(\Omega)}$ (11)
$\hskip 56.9055pt\leq\|y_{\varrho}-y_{h}\|_{L^{2}(\Gamma)}^{2}+\varrho\,\|y_{\varrho}-y_{h}\|^{2}_{H^{1}(\Omega)}\quad\mbox{for all}\;y_{h}\in Y_{h}.$
Lemma 3.1.
Let $y_{\varrho h}\in Y_{h}$ be the unique solution of (10). Then, for $\overline{y}\in L^{2}(\Gamma)$, there holds the error estimate
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq(1+\sqrt{2})\,\|\overline{y}\|_{L^{2}(\Gamma)},$ (12)
while, for $\overline{y}\in H^{1/2}(\Gamma)$ and choosing $\varrho=h$, we have
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq c\,h^{1/2}\,\|\overline{y}\|_{H^{1/2}(\Gamma)}.$ (13)
A related result holds for $\overline{y}\in H^{1}(\Gamma)$ where, for $\varrho=h$,
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq c\,h\,\|\overline{y}\|_{H^{1}(\Gamma)}.$ (14)
Proof.
For $\overline{y}\in L^{2}(\Gamma)$, we consider (11) for $y_{h}=0$, and we use (5) and (6) to obtain
$\|y_{\varrho}-y_{\varrho h}\|_{L^{2}(\Gamma)}^{2}\leq\|y_{\varrho}\|_{L^{2}(\Gamma)}^{2}+\varrho\,\|y_{\varrho}\|^{2}_{H^{1}(\Omega)}\leq 2\,\|\overline{y}\|_{L^{2}(\Gamma)}^{2}.$
With the triangle inequality and again using (6) we therefore conclude
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq\|y_{\varrho h}-y_{\varrho}\|_{L^{2}(\Gamma)}+\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}\leq(1+\sqrt{2})\,\|\overline{y}\|_{L^{2}(\Gamma)}\,.$
Next we consider the case when $\overline{y}_{e}\in Y$ is the extension of $\overline{y}$ as solution of the biharmonic boundary value problem (9). When using the triangle inequality twice in combination with Cea’s estimate (11), we can write, for arbitrary $y_{h}\in Y_{h}$
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}^{2}$ $\leq$ $2\,\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}^{2}+2\,\|y_{\varrho}-y_{\varrho h}\|^{2}_{L^{2}(\Gamma)}$
$\leq$ $2\,\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}^{2}+2\,\|y_{\varrho}-y_{h}\|_{L^{2}(\Gamma)}^{2}+2\,\varrho\,\|y_{\varrho}-y_{h}\|^{2}_{H^{1}(\Omega)}$
$\leq$ $6\,\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}^{2}+4\,\|\overline{y}-y_{h}\|_{L^{2}(\Gamma)}^{2}$
+ $\hskip 56.9055pt+\,4\,\varrho\,\|y_{\varrho}-\overline{y}_{e}\|_{H^{1}(\Omega)}^{2}+4\,\varrho\,\|\overline{y}_{e}-y_{h}\|^{2}_{H^{1}(\Omega)}.$
In particular for $y_{h}=P_{h}\overline{y}_{e}\in Y_{h}$ being the $L^{2}$ projection of $\overline{y}_{e}$, we have the standard fe error estimates
$\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{H^{1}(\Omega)}\leq c\,\|\overline{y}_{e}\|_{H^{1}(\Omega)},\quad\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{L^{2}(\Omega)}\leq c\,h\,\|\overline{y}_{e}\|_{H^{1}(\Omega)}.$ (16)
When using [1, Theorem 3.6] and a space interpolation argument, we also have
$\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{L^{2}(\Gamma)}\leq c\,\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{H^{1/2}(\Omega)}\leq c\,h^{1/2}\,\|\overline{y}_{e}\|_{H^{1}(\Omega)}.$
Hence, using (7) and $\|\overline{y}_{e}\|_{H^{1}(\Omega)}\leq\|\overline{y}\|_{H^{1/2}(\Gamma)}$, this gives
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}^{2}\,\leq\,c\,(\varrho+h)\,\|\overline{y}\|_{H^{1/2}(\Gamma)}^{2},$
and (13) follows when choosing $\varrho=h$.
Finally, we consider the case $\overline{y}\in H^{1}(\Gamma)$ which implies $\overline{y}_{e}\in H^{3/2}(\Omega)$. In this case, we have
$\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{H^{1}(\Omega)}\leq c\,h^{1/2}\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)},\;\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{L^{2}(\Omega)}\leq c\,h^{3/2}\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)},$
and
$\|\overline{y}-P_{h}\overline{y}_{e}\|_{L^{2}(\Gamma)}\leq c\,h\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)}.$
Together with (8) we therefore conclude
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}^{2}\,\leq\,c\,(\varrho^{2}+\varrho h+h^{2})\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)}^{2},$
and for $\varrho=h$ this gives (14). ∎
While the regularization error estimates as given in Lemma 2.1 are optimal in $\varrho$ for $\overline{y}\in H^{1}(\Gamma)$, i.e., $\overline{y}_{e}\in H^{3/2}(\Omega)$, we can expect higher order convergence for the fe approximation $P_{h}\overline{y}_{e}$ when $\overline{y}_{e}$ is more regular. In particular for $\overline{y}_{e}\in H^{2}(\Omega)$ we have
$\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{H^{1}(\Omega)}\leq c\,h\,\|\overline{y}_{e}\|_{H^{2}(\Omega)},\;\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{L^{2}(\Omega)}\leq c\,h^{2}\,\|\overline{y}_{e}\|_{H^{2}(\Omega)}.$
Since $\overline{y}_{e}$ is defined as solution of the biharmonic boundary value problem (9), we have to formulate additional compatibility conditions in order to ensure $\overline{y}_{e}\in H^{2}(\Omega)$, when $\Gamma=\partial\Omega$ is piecewise smooth. In particular, the tangential derivatives of $\overline{y}$ have to be in $H^{1/2}(\Gamma)$, see, e.g., the discussion in [2, 6] in the two-dimensional case. Then we can use (3) to conclude
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}^{2}\,\leq\,c_{1}\,\varrho^{2}\,\|\overline{y}\|_{H^{1}(\Gamma)}^{2}+c_{2}\,(h^{3}+\varrho\,h^{2})\,\|\overline{y}_{e}\|^{2}_{H^{2}(\Omega)},$
and for $\varrho\leq h^{3/2}$ we obtain
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\,\leq\,c\,h^{3/2}\,\Big[\|\overline{y}_{e}\|^{2}_{H^{2}(\Omega)}+\|\overline{y}\|_{H^{3/2}(\Gamma)}^{2}\Big]^{1/2}.$ (17)
However, for $\overline{y}\in H^{2}(\Gamma)$, i.e., $\overline{y}_{e}\in H^{5/2}(\Omega)$, and using the best approximation estimate for the boundary term, this gives
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}^{2}\,\leq\,c_{1}\,h^{4}\,\|\overline{y}\|_{H^{2}(\Gamma)}^{2}+c_{2}\,\varrho^{2}\,\|\overline{y}\|_{H^{1}(\Gamma)}^{2}+c_{3}\,\varrho\,h^{2}\,\|\overline{y}_{e}\|^{2}_{H^{2}(\Omega)},$
and for $\varrho=h^{2}$ we finally obtain
$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\,\leq\,c\,h^{2}\Big[\|\overline{y}\|_{H^{2}(\Gamma)}^{2}+\|\overline{y}\|_{H^{1}(\Gamma)}^{2}+\|\overline{y}_{e}\|^{2}_{H^{2}(\Omega)}\Big]^{1/2}.$ (18)
4 Fast solvers
Once the basis is chosen, the finite element scheme (10) is equivalent to a linear system of finite element equations that can be written in the form
$\widetilde{\mathbf{M}}_{h}+\varrho(\widetilde{\mathbf{K}}_{h}+\mathring{\mathbf{K}}_{h})]{\mathbf{y}}_{h}=\mathbf{\overline{y}}_{h},$ (19)
where the matrices $\widetilde{\mathbf{M}}_{h}$, $\widetilde{\mathbf{K}}_{h}$, and $\mathring{\mathbf{K}}_{h}$ have the respective block representations
$\widetilde{\mathbf{M}}_{h}=\left(\begin{array}[]{cc}{\mathbf{0}}&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{M}}_{BB}\end{array}\right),\quad\widetilde{\mathbf{K}}_{h}=\left(\begin{array}[]{cc}{\mathbf{0}}&{\mathbf{0}}\\ {\mathbf{0}}&\widetilde{\mathbf{K}}_{BB}\end{array}\right),\quad\mbox{and}\quad\mathring{\mathbf{K}}_{h}=\left(\begin{array}[]{cc}\mathring{\mathbf{K}}_{II}&\mathring{\mathbf{K}}_{IB}\\ \mathring{\mathbf{K}}_{BI}&\mathring{\mathbf{K}}_{BB}\end{array}\right),$
when we split the unknowns (dofs) $\mathbf{y}_{h}=(\mathbf{y}_{I}^{\top},\mathbf{y}_{B}^{\top})^{\top}\in\mathbb{R}^{(n-1)^{d}}$ into strict interior unknowns $\mathbf{y}_{I}\in\mathbb{R}^{(n-3)^{d}}$ and near-boundary unknowns $\mathbf{y}_{B}\in\mathbb{R}^{(n-1)^{d}-(n-3)^{d}}$. The matrices ${\mathbf{M}}_{BB}$, $\widetilde{\mathbf{K}}_{BB}$ and $\mathring{\mathbf{K}}_{h}$ are defined by the identities $({\mathbf{M}}_{BB}\mathbf{y}_{B},\mathbf{v}_{B})=\langle y_{h},v_{h}\rangle_{L^{2}(\Gamma)}$, $(\widetilde{\mathbf{K}}_{BB}\mathbf{y}_{B},\mathbf{v}_{B})=\langle y_{h},v_{h}\rangle_{H^{1}(\widetilde{\Omega}_{h})}$, $(\mathring{\mathbf{K}}_{h}\mathbf{y}_{h},\mathbf{v}_{h})=\langle y_{h},v_{h}\rangle_{H^{1}(\mathring{\Omega}_{h})}$ for all $\mathbf{y}_{h}=(\mathbf{y}_{I}^{\top},\mathbf{y}_{B}^{\top})^{\top}\leftrightarrow y_{h},v_{h}\in Y_{h}$ (fe isomorphism), where $\mathring{\Omega}_{h}=\Omega\setminus\overline{\widetilde{\Omega}}_{h}=(h,1-h)^{d}$ and $\widetilde{\Omega}_{h}=\Omega\setminus\overline{\mathring{\Omega}}_{h}$, whereas $\mathbf{\overline{y}}_{h}=(\mathbf{0}_{I}^{\top},\mathbf{\overline{y}}_{B}^{\top})^{\top}\in\mathbb{R}^{m}$ is given by $(\mathbf{\overline{y}}_{B},\mathbf{y}_{B})=\langle\overline{y}_{h},y_{h}\rangle_{L^{2}(\Gamma)}$ for all $\mathbf{y}_{h}=(\mathbf{y}_{I}^{\top},\mathbf{y}_{B}^{\top})^{\top}\leftrightarrow y_{h},v_{h}\in Y_{h}$.
Eliminating $\mathbf{y}_{I}=-\mathring{\mathbf{K}}_{II}^{-1}\mathring{\mathbf{K}}_{IB}\mathbf{y}_{B}$ from the linear system (19), we arrive at the boundary Schur complement system
${\mathbf{S}}_{BB}\mathbf{y}_{B}=\mathbf{\overline{y}}_{B}$ (20)
with ${\mathbf{S}}_{BB}={\mathbf{M}}_{BB}+\varrho(\widetilde{\mathbf{K}}_{BB}+\mathring{\mathbf{S}}_{BB})={\mathbf{M}}_{BB}+\varrho(\widetilde{\mathbf{K}}_{BB}+(\mathring{\mathbf{K}}_{BB}-\mathring{\mathbf{K}}_{BI}\mathring{\mathbf{K}}_{II}^{-1}\mathring{\mathbf{K}}_{IB}))$. The Schur complement system (20) can efficiently be solved by means of the Conjugate Gradient (CG) method without any preconditioning since, for $\varrho\leq h$, the Schur complement ${\mathbf{S}}_{BB}$ is spectrally equivalent to the boundary mass matrix $\mathbf{M}_{BB}$, and in turn $\mathbf{M}_{BB}$ is spectrally equivalent to the lumped boundary mass matrix $\mbox{lump}(\mathbf{M}_{BB})$ and to $h^{d-1}{\mathbf{I}}_{BB}$. Indeed, it is easy to see that
${\mathbf{M}}_{BB}\leq{\mathbf{S}}_{BB}={\mathbf{M}}_{BB}+\varrho(\widetilde{\mathbf{K}}_{BB}+\mathring{\mathbf{S}}_{BB})\leq(1+\widetilde{c}\varrho h^{-1}+\mathring{c}\varrho h^{-1}){\mathbf{M}}_{BB}$ (21)
with $h$ and $\varrho$ independent positive constants $\widetilde{c}$ and $\mathring{c}$. The choice $\varrho\leq h$ delivers the desired result. It is recommended to use $\mbox{lump}(\mathbf{M}_{BB})$ as diagonal preconditioner in the Preconditioned Conjugate Gradient (PCG) method since it provides the right scaling. The numerical results presented in Section 5 show that the system (19) can also efficiently be solved by means of PCG with a simple Algebraic MultiGrid (AMG) preconditioner.
5 Numerical results
We first consider the target
$\overline{y}=\overline{y}(x):=\cos(\pi x_{1})\cos(\pi x_{2})\cos(\pi x_{3}),\;x=(x_{1},x_{2},x_{3})\in\Gamma,$ (22)
on the boundary $\Gamma=\partial\Omega$ of the domain $\Omega=(0,1)^{3}$. We mention that $\overline{y}$ is the trace of a smooth function with vanishing normal derivative on the boundary $\Gamma$, i.e., we have $\overline{y}\in H^{2}(\Gamma)$, and the error estimate (18) applies when choosing $\varrho=h^{2}$. We use a tensor product mesh as described in Section 3. The initial mesh contains $5$ vertices in each direction, and $125$ in total with mesh size $h=0.25$. We note that we have only $3$ dofs in each direction, and $27$ in total for the initial level. Table 1 provides the numerical results starting from level $\ell=1$ with $27$ dofs and running to the finest discretization level $\ell=7$ obtained by $6$ uniform refinements of the initial mesh. The fourth column displays the $L^{2}$ error $\|y-\overline{y}\|_{L^{2}(\Gamma)}$ on the boundary. As expected, we observe second order of convergence; cf. experimental order of convergence (eoc) given in the fifth column. We first solve the original system (19) by means of AMG preconditioned CG iterations (#AMG-PCG its), and observe that not more than $4$ iterations are needed in order to reach a relative residual error of $10^{-9}$. We further test the CG and lumped mass preconditioned CG solvers for the Schur complement equation (20) until the relative residual error reaches $10^{-9}$. The number of Schur complement CG (#SCG its) and lumped mass preconditioned CG (#SPCG its) iterations are displayed in the last two columns of Table 1. As expected from the theoretical results given in Section 4, we see level-independent iteration numbers in both cases. Moreover, the lumped-mass preconditioner further reduces the number of iteration by the scaling effect. We note that the action of $\mathring{\mathbf{K}}_{II}^{-1}$ to a vector within the multiplication of the Schur complement $\mathbf{S}_{BB}$ by some vector (iterate) is realized by an AMG preconditioned CG method until the relative residual error is reduced by a factor $10^{10}$. The latter accuracy of this inner PCG iteration can be adapted (reduced !) to the outer CG/PCG iteration following the results from [7].
$\ell$ #dofs $h$ error eoc #AMG-PCG its #SCG its #PSCG its
11 $27$ $2^{-2}$ 1.6691.669e1-1 - 22 11 11
22 $343$ $2^{-3}$ 5.2155.215e2-2 $1.68$ 33 66 66
33 $3,375$ $2^{-4}$ 1.3471.347e2-2 $1.95$ 44 $17$ 99
44 $29,791$ $2^{-5}$ 3.2263.226e3-3 $2.06$ 44 $24$ 99
55 $250,047$ $2^{-6}$ 7.7077.707e4-4 $2.07$ 44 $28$ 88
66 $2,048,383$ $2^{-7}$ 1.8721.872e4-4 $2.04$ 44 $29$ 66
77 $16,581,375$ $2^{-8}$ 4.6054.605e5-5 $2.02$ 44 $29$ 44
Table 1: Target (22): error = $\|y-\overline{y}\|_{L^{2}(\Gamma)}$, number of AMG CG iterations (#ACG-PCG its) for the original system (19), and number of CG (#SCG its) and lumped-mass preconditioned CG (#PSCG its) iterations for the Schur complement system (20), $\varrho=h^{2}$.
The second target
$\overline{y}(x):=x_{1}^{2}-0.5x_{2}^{2}-0.5x_{3}^{2},\;x=(x_{1},x_{2},x_{3})\in\Gamma,$ (23)
is a trace of a smooth function which does not fulfill the homogeneous Neumann boundary conditions on the boundary $\Gamma$ of $\Omega=(0,1)^{3}$, i.e., $\overline{y}\in H^{3/2-\varepsilon}(\Gamma)$, and $\overline{y}_{e}\in H^{2-\varepsilon}(\Omega)$, $\varepsilon>0$. Similar as in (17) we therefore expect a reduced eco of about 1.5. We perform the same tests for the target (23) as for the previous example for the target (22). The results are given in Table 2, where we have used $\varrho=h^{3/2}$ as prescribed by the theory. The iteration numbers are again independent of $h$, and show the same behavior as in the case of the first example.
$\ell$ #Dofs $h$ error eoc #AMG-PCG its #SCG its #PSCG its
11 $27$ $2^{-2}$ 4.8694.869e1-1 - 22 55 55
22 $343$ $2^{-3}$ 2.1312.131e1-1 $1.19$ 33 $13$ $10$
33 $3,375$ $2^{-4}$ 8.1778.177e2-2 $1.38$ 55 $14$ 88
44 $29,791$ $2^{-5}$ 2.9462.946e2-2 $1.47$ 44 $15$ 88
55 $250,047$ $2^{-6}$ 1.0361.036e2-2 $1.51$ 44 $16$ 77
66 $2,048,383$ $2^{-7}$ 3.6193.619e3-3 $1.52$ 44 $18$ 77
77 $16,581,375$ $2^{-8}$ 1.2651.265e3-3 $1.52$ 44 1919 77
Table 2: Target (23): Same agenda as Table 1, but now $\varrho=h^{3/2}$.
6 Conclusions and Outlook
We have investigated a boundary value tracking, distributed ocp for some elliptic bvp with homogeneous Neumann boundary conditions. This ocp can be reduced to a state-based variational problem for functions from $H^{1}(\Omega)$ with vanishing normal derivatives in $H^{-1/2}(\Gamma)$. We have constructed a conforming fe discretization for the special case of tensor-product meshes, and we have derived discretization error estimates and fast solvers. The numerical experiments illustrate the theoretical results quantitatively. In general, one has to include the homogeneous Neumann boundary conditions by means of Lagrange multipliers, which is a topic of future research.
Acknowledgement
We would like to thank RICAM for the computing resource support. The financial support for the fourth author by the Austrian Federal Ministry for Digital and Economic Affairs, the National Foundation for Research, Technology and Development, and the Christian Doppler Research Association is gratefully acknowledged.
Appendix
In this appendix we provide the proofs of all results as used in this paper, where $\widetilde{S}_{h}^{1}(0,1)$ is the modified finite element space of piecewise linear and continuous basis functions $\varphi_{i}(x)$ for $x\in(0,1)$ with zero derivatives $\varphi_{i}^{\prime}(0)=\varphi_{i}^{\prime}(1)=0$.
Lemma 6.1.
Let $y\in C([0,1])$ be a given continuous function with $y^{\prime}(0)=y^{\prime}(1)=0$, and satisfying $y^{\prime\prime}\in L^{2}(0,1)$. For the piecewise linear interpolation
$I_{h}y(x)=\sum\limits_{i=1}^{n-1}y(x_{i})\varphi_{i}(x).$
there hold the error estimates
$\|y-I_{h}y\|_{L^{2}(0,1)}\leq\frac{1}{2}\,h^{2}\,\|y^{\prime\prime}\|_{L^{2}(0,1)},$ (24)
$\|y-I_{h}y\|_{L^{2}(0,1)}\leq\sqrt{2}\,h\,\|y^{\prime}\|_{L^{2}(0,1)},$ (25)
and
$\|y^{\prime}-(I_{h}y)^{\prime}\|_{L^{2}(0,1)}\leq\frac{1}{\sqrt{2}}\,h\,\|y^{\prime\prime}\|_{L^{2}(0,1)}.$ (26)
Moreover, the interpolation operator is stable in $H^{1}(0,1)$, i.e.,
$\|(I_{h}y)^{\prime}\|_{L^{2}(0,1)}\leq\|y^{\prime}\|_{L^{2}(0,1)}\quad\mbox{for all}\;y\in H^{1}(0,1).$ (27)
Proof.
For $i=2,\ldots,n-1$ we have the standard interpolation error estimates
$\int_{x_{i-1}}^{x_{i}}[y(x)-I_{h}y(x)]^{2}\,dx\leq\frac{1}{24}\,h^{4}\,\int_{x_{i-1}}^{x_{i}}[y^{\prime\prime}(x)]^{2}\,dx\,$,
and
$\int_{x_{i-1}}^{x_{i}}[y^{\prime}(x)-(I_{h}y)^{\prime}(x)]^{2}\,dx\leq\frac{1}{3}\,h^{2}\,\int_{x_{i-1}}^{x_{i}}[y^{\prime\prime}(x)]^{2}\,dx\,$,
For the first interval $(x_{0},x_{1})$ we have $I_{h}y(x_{1})=y(x_{1})$, and hence we can write for $x\in(x_{0},x_{1})$
$y(x)-I_{h}y(x)=[I_{h}y(x_{1})-I_{h}y(x)]-[y(x_{1})-y(x)]=\int_{x}^{x_{1}}[(I_{h}y)^{\prime}(s)-y^{\prime}(s)]\,ds,$
and with the Cauchy–Schwarz inequality we obtain
$[y(x)-I_{h}y(x)]^{2}$ $=$ $\left(\int_{x}^{x_{1}}[(I_{h}y)^{\prime}(s)-y^{\prime}(s)]\,ds\right)^{2}$
$\leq$ $\int_{x}^{x_{1}}1^{2}\,ds\int_{x}^{x_{1}}[(I_{h}y)^{\prime}(s)-y^{\prime}(s)]^{2}\,ds$
$\leq$ $(x_{1}-x)\int_{x_{0}}^{x_{1}}[(I_{h}y)^{\prime}(s)-y^{\prime}(s)]^{2}\,ds,$
i.e.,
$\int_{x_{0}}^{x_{1}}[y(x)-I_{h}y(x)]^{2}\,dx\leq\frac{1}{2}\,h^{2}\,\int_{x_{0}}^{x_{1}}[(I_{h}y)^{\prime}(x)-y^{\prime}(x)]^{2}\,dx.$
Due to $y^{\prime}(x_{0})=(I_{h}y)^{\prime}(x_{0})=0$ we further have
$(I_{h}y)^{\prime}(x)-y^{\prime}(x)$ $=$ $[(I_{h}y)^{\prime}(x)-(I_{h}y)^{\prime}(x_{0})]-[y^{\prime}(x)-y^{\prime}(x_{0})]$
$=$ $\int_{x_{0}}^{x}[(I_{h}y)^{\prime\prime}(s)-y^{\prime\prime}(s)]\,ds\,=\,\int_{x_{0}}^{x}[-y^{\prime\prime}(s)]\,ds,$
and hence,
$[(I_{h}y)^{\prime}(x)-y^{\prime}(x)]^{2}$ $=$ $\left(\int_{x_{0}}^{x}[-y^{\prime\prime}(s)]\,ds\right)^{2}\leq\int_{x_{0}}^{x}1^{2}\,ds\int_{x_{0}}^{x}[y^{\prime\prime}(s)]^{2}\,ds$
$\leq$ $(x-x_{0})\int_{x_{0}}^{x_{1}}[y^{\prime\prime}(s)]^{2}\,ds,$
i.e.,
$\int_{x_{0}}^{x_{1}}[(I_{h}y)^{\prime}(x)-y^{\prime}(x)]^{2}dx\leq\frac{1}{2}\,h^{2}\,\int_{x_{0}}^{x_{1}}[y^{\prime\prime}(x)]^{2}\,dx$
follows, and we also conclude
$\int_{x_{0}}^{x_{1}}[y(x)-I_{h}y(x)]^{2}\,dx\leq\frac{1}{4}\,h^{4}\,\int_{x_{0}}^{x_{1}}[y^{\prime\prime}(x)]^{2}\,dx.$
For the last interval $(x_{n-1},x_{n})$ the proof follows the same lines. When summing up the local contributions, the error estimates (24) and (26) follow. Moreover, for $i=2,\ldots,n-1$ and $x\in(x_{i-1},x_{i})$, we write
$(I_{h}y)^{\prime}(x)=\frac{1}{h}[y(x_{i})-y(x_{i-1})]=\frac{1}{h}\int_{x_{i-1}}^{x_{i}}y^{\prime}(s)\,ds,$
i.e.,
$(I_{h}y)^{\prime}(x)]^{2}=\left(\frac{1}{h}\int_{x_{i-1}}^{x_{i}}y^{\prime}(s)\,ds\right)^{2}\leq\frac{1}{h}\int_{x_{i-1}}^{x_{i}}[y^{\prime}(s)]^{2}\,ds.$
Integration and summing up over all $i=2,\ldots,n-1$ gives, recall $(I_{h}y)^{\prime}(x)=0$ for $x\in[x_{0},x_{1}]\cup[x_{n-1},x_{n}]$,
$\|(I_{h}y)^{\prime}\|^{2}_{L^{2}(0,1)}=\sum\limits_{i=2}^{n-1}\|(I_{h}y)^{\prime}\|^{2}_{L^{2}(x_{i-1},x_{i})}\leq\sum\limits_{i=2}^{n-1}\|y^{\prime}\|^{2}_{L^{2}(x_{i-1},x_{i})}\leq\|y^{\prime}\|^{2}_{L^{2}(0,1)},$
i.e., (27). With this we finally conclude (25),
$\|y-I_{h}y\|_{L^{2}(0,1)}\leq\frac{1}{\sqrt{2}}\,h\,\|y^{\prime}-(I_{h}y)^{\prime}\|_{L^{2}(0,1)}\leq\sqrt{2}\,h\,\|y^{\prime}\|_{L^{2}(0,1)}.$
Next we introduce the $L^{2}$ projection $Q_{h}:L^{2}(0,1)\to\widetilde{S}^{1}_{h}(0,1)$, and present related error and stability estimates.
Lemma 6.2.
For any given $y\in L^{2}(0,1)$ we define $Q_{h}y\in\widetilde{S}_{h}^{1}(0,1)$ as unique solution of the variational formulation
$\langle Q_{h}y,y_{h}\rangle_{L^{2}(0,1)}=\langle y,y_{h}\rangle_{L^{2}(0,1)}\quad\mbox{for all}\;y_{h}\in\widetilde{S}_{h}^{1}(0,1).$ (28)
Then,
$\|Q_{h}y\|_{L^{2}(0,1)}\leq\|y\|_{L^{2}(0,1)}\quad\mbox{for all}\;y\in L^{2}(0,1).$ (29)
For $y\in H^{2}(0,1)$ with $y^{\prime}(0)=y^{\prime}(1)=0$ there hold the error estimates
$\|y-Q_{h}y\|_{L^{2}(0,1)}\leq\frac{1}{2}\,h^{2}\,\|y^{\prime\prime}\|_{L^{2}(0,1)},$ (30)
and
$\|y^{\prime}-(Q_{h}y)^{\prime}\|_{L^{2}(0,1)}\leq c\,h\,\|y^{\prime\prime}\|_{L^{2}(0,1)}.$ (31)
Moreover, there holds the stability estimate
$\|(Q_{h}y)^{\prime}\|_{L^{2}(0,1)}\leq c\,\|y^{\prime}\|_{L^{2}(0,1)}\quad\mbox{for all}\;y\in H^{1}(0,1),y^{\prime}(0)=y^{\prime}(1)=0.$ (32)
Proof.
The stability estimate (29) is a consequence of the variational formulation (28) when choosing $y_{h}=Q_{h}y$. By Cea’s lemma and using (24) we have (30),
$\|y-Q_{h}y\|_{L^{2}(0,1)}\leq\inf\limits_{y_{h}\in Y_{h}}\|y-y_{h}\|_{L^{2}(0,1)}\leq\|y-I_{h}y\|_{L^{2}(0,1)}\leq\frac{1}{2}\,h^{2}\,\|y^{\prime\prime}\|_{L^{2}(0,1)}.$
Moreover, using the triangle inequatlity twice, the inverse inequality
$\|y_{h}^{\prime}\|_{L^{2}(0,1)}\leq\sqrt{12}\,h^{-1}\,\|y_{h}\|_{L^{2}(0,1)}\quad\mbox{for}\;y_{h}\in\widetilde{S}_{h}^{1}(0,1),$
and the error extimates (26) as well as (24) and (30), this gives (31),
$\|y^{\prime}-(Q_{h}y)^{\prime}\|_{L^{2}(0,1)}$ $\leq$ $\|y^{\prime}-(I_{h}y)^{\prime}\|_{L^{2}(0,1)}+\|(I_{h}y)^{\prime}-(Q_{h}y)^{\prime}\|_{L^{2}(0,1)}$
$\hskip-71.13188pt\leq\,\frac{1}{\sqrt{2}}\,h\,\|y^{\prime\prime}\|_{L^{2}(0,1)}+2\sqrt{3}\,h^{-1}\,\|I_{h}y-Q_{h}y\|_{L^{2}(0,1)}$
$\hskip-71.13188pt\leq\,\frac{1}{\sqrt{2}}\,h\,\|y^{\prime\prime}\|_{L^{2}(0,1)}+2\sqrt{3}\,h^{-1}\,\Big[\|I_{h}y-y\|_{L^{2}(0,1)}+\|y-Q_{h}y\|_{L^{2}(0,1)}\Big]$
$\hskip-71.13188pt\leq\,\frac{1}{2}\Big(\sqrt{2}+4\sqrt{3}\Big)\,h\,\|y^{\prime\prime}\|_{L^{2}(0,1)}.$
When using the same arguments we finally have
$\|(Q_{h}y)^{\prime}\|_{L^{2}(0,1)}$ $\leq$ $\|(Q_{h}y)^{\prime}-(I_{h}y)^{\prime}\|_{L^{2}(0,1)}+\|(I_{h}y)^{\prime}\|_{L^{2}(0,1)}$
$\leq$ $2\sqrt{3}\,h^{-1}\|Q_{h}y-I_{h}y\|_{L^{2}(0,1)}+\|y^{\prime}\|_{L^{2}(0,1)}$
$\leq$ $2\sqrt{3}\,h^{-1}\Big[\|Q_{h}y-y\|_{L^{2}(0,1)}+\|y-I_{h}y\|_{L^{2}(0,1)}\Big]+\|y^{\prime}\|_{L^{2}(0,1)}$
$\leq$ $(1+4\sqrt{6})\,\|y^{\prime}\|_{L^{2}(0,1)}.$
For $\Omega=(0,1)^{d}$ we use the tensor product finite element space $Y_{h}=\otimes_{i=1}^{d}\widetilde{S}_{h}^{1}(0,1)$, and we define $P_{h}:=\otimes_{i=1}^{d}Q_{h_{i}}:Y\to Y_{h}$, where $Q_{h_{i}}:L^{2}(0,1)\to\widetilde{S}_{h}^{1}(0,1)$ are defined as in (28), but with respect to the component $x_{i}\in{\mathbb{R}}$ of $x\in{\mathbb{R}}^{d}$.
Lemma 6.3.
For $y\in Y\cap H^{2}(\Omega)$ and $P_{h}y\in Y_{h}$, there hold the error estimates
$\|y-P_{h}y\|_{L^{2}(\Omega)}\leq c\,h^{2}\,|y|_{H^{2}(\Omega)},$ (33)
and
$\|\nabla(y-P_{h}y)\|_{L^{2}(\Omega)}\leq c\,h\,|y|_{H^{2}(\Omega)}.$ (34)
Proof.
For simplity, we consider the case $d=2$ only. Then, we conclude
$\|y-P_{h}y\|_{L^{2}(\Omega)}$ $=$ $\|y-Q_{h_{1}}Q_{h_{2}}y\|_{L^{2}(\Omega)}$
$\leq$ $\|y-Q_{h_{1}}y\|_{L^{2}(\Omega)}+\|Q_{h_{1}}(y-Q_{h_{2}}y)\|_{L^{2}(\Omega)}$
$\leq$ $\|y-Q_{h_{1}}y\|_{L^{2}(\Omega)}+c\,\|y-Q_{h_{2}}y\|_{L^{2}(\Omega)}$
$\leq$ $c_{1}\,h_{1}^{2}\,\|\partial^{2}_{y_{1}y_{1}}y\|_{L^{2}(\Omega)}+c_{2}\,h_{2}^{2}\,\|\partial^{2}_{y_{2}y_{2}}y\|_{L^{2}(\Omega)}$
$\leq$ $c\,h^{2}\,|y|_{H^{2}(\Omega)},$
when using the stability estimate (29), and the error estimate (30), i.e., (33). The proof of (34) follows the same lines, but using the stability estimate (32), and the error estimate (31). ∎
It remains to prove a best approximation result for the boundary term as used to derive the error estimate (18).
Lemma 6.4.
Let $\Gamma=\partial\Omega$ for $\Omega=(0,1)^{d}$. Assume $y\in H^{2}(\Gamma)$, i.e., there exists an extension $y_{e}\in Y\cap H^{5/2}(\Omega)$ as solution of the biharmonic boundary value problem (9). Then, there holds the error estimate
$\|y-P_{h}y\|_{L^{2}(\Gamma)}\leq c\,h^{2}\,|y|_{H^{2}(\Gamma)}.$ (35)
Proof.
Again, we consider the case $d=2$ only. Let $\Gamma_{0}:=\{(x,0):x\in(0,1)\}$. With (24) and (29) we then have
$\|y-P_{h}y\|_{L^{2}(\Gamma_{0})}$ $=$ $\|y-Q_{h_{1}}Q_{h_{2}}y\|_{L^{2}(\Gamma_{0})}$
$\leq$ $\|y-Q_{h_{1}}y\|_{L^{2}(\Gamma_{0})}+\|Q_{h_{1}}(y-Q_{h_{2}}y)\|_{L^{2}(\Gamma_{0})}$
$\leq$ $\frac{1}{2}\,h^{2}\,\|\partial_{x_{1}x_{1}}y\|_{L^{2}(\Gamma_{0})}+\|y-Q_{h_{2}}y\|_{L^{2}(\Gamma_{0})}$
$\leq$ $\frac{1}{2}\,h^{2}\,\|y\|_{H^{5/2}(\Omega)}+\|y-I_{h_{2}}y\|_{L^{2}(\Gamma_{0})}.$
For $x\in\Gamma_{0}$ we can write, using $\partial_{x_{2}}y_{e}(x,s)_{|s=0}=0$,
$y(x,0)-I_{h_{2}}y(x,0)$ $=$ $y_{e}(x,0)-y_{e}(x,h)=-\int_{0}^{h}\partial_{x_{2}}y_{e}(x,s)\,ds$
$\hskip-56.9055pt=\,\int_{0}^{h}[\partial_{x_{2}}y_{e}(x,0)-\partial_{x_{2}}y_{e}(x,s)]\,ds\,=\,-\int_{0}^{h}\int_{0}^{s}\partial_{x_{2}x_{2}}y_{e}(x,\tau)\,d\tau ds.$
Hence we obtain
$[y(x,0)-I_{h_{2}}y(x,0)]^{2}$ $=$ $\left[\int_{0}^{h}\int_{0}^{s}1\cdot\partial_{x_{2}x_{2}}y_{e}(x,\tau)\,d\tau ds\right]^{2}$
$\leq$ $\int_{0}^{h}1^{2}\,ds\int_{0}^{h}\left[\int_{0}^{s}1\cdot\partial_{x_{2}x_{2}}y_{e}(x,\tau)\,d\tau\right]^{2}ds$
$\leq$ $h\int_{0}^{h}\int_{0}^{s}1^{2}\,d\tau\int_{0}^{s}[\partial_{x_{2}x_{2}}y_{e}(x,\tau)]^{2}d\tau ds$
$\leq$ $h\int_{0}^{h}s\,ds\int_{0}^{h}[\partial_{x_{2}x_{2}}y_{e}(x,\tau)]^{2}d\tau$
$=$ $\frac{1}{2}\,h^{3}\,\int_{0}^{h}[\partial_{x_{2}x_{2}}y_{e}(x,\tau)]^{2}d\tau.$
Next we consider $x\in(x_{i-1},x_{i})$, $i=1,\ldots,n$, and use the piecewise constant $L^{2}$ projection
$\hat{y}_{i}(\tau)=\frac{1}{h}\int_{x_{i-1}}^{x_{i}}\partial_{x_{2}x_{2}}y_{e}(\eta,\tau)\,d\eta,$
to write
$\partial_{x_{2}x_{2}}y_{e}(x,\tau)]^{2}\leq 2\,[\partial_{x_{2}x_{2}}y_{e}(x,\tau)-\hat{y}_{i}(\tau)]^{2}+2\,[\hat{y}_{i}(\tau)]^{2},$
i.e.,
$y(x,0)-I_{h_{2}}y(x,0)]^{2}\,\leq\,h^{3}\int_{0}^{h}[\partial_{x_{2}x_{2}}y_{e}(x,\tau)-\hat{y}_{i}(\tau)]^{2}d\tau+h^{3}\int_{0}^{h}[\hat{y}_{i}(\tau)]^{2}d\tau,$
and integration over $x\in(x_{i-1},x_{i})$ gives
$\int_{x_{i-1}}^{x_{i}}[y(x,0)-I_{h_{2}}y(x,0)]^{2}dx$
$\hskip 28.45274pt\leq\,h^{3}\int_{x_{i-1}}^{x_{i}}\int_{0}^{h}[\partial_{x_{2}x_{2}}y_{e}(x,\tau)-\hat{y}_{i}(\tau)]^{2}d\tau dx+h^{3}\int_{x_{i-1}}^{x_{i}}\int_{0}^{h}[\hat{y}_{i}(\tau)]^{2}d\tau dx.$
Using
$\int_{x_{i-1}}^{x_{i}}[\partial_{x_{2}x_{2}}y_{e}(x,\tau)-\hat{y}_{i}(\tau)]^{2}dx$ $=$ $\int_{x_{i-1}}^{x_{i}}\left[\partial_{x_{2}x_{2}}y_{e}(x,\tau)-\frac{1}{h}\int_{x_{i-1}}^{x_{i}}\partial_{x_{2}x_{2}}y_{e}(\eta,\tau)\,d\eta\right]^{2}dx$
$\hskip-85.35826pt=\,\int_{x_{i-1}}^{x_{i}}\left[\frac{1}{h}\int_{x_{i-1}}^{x_{i}}[\partial_{x_{2}x_{2}}y_{e}(x,\tau)-\partial_{x_{2}x_{2}}y_{e}(\eta,\tau)]\,d\eta\right]^{2}dx$
$\hskip-85.35826pt=\,\frac{1}{h^{2}}\int_{x_{i-1}}^{x_{i}}\left[\int_{x_{i-1}}^{x_{i}}\frac{\partial_{x_{2}x_{2}}y_{e}(x,\tau)-\partial_{x_{2}x_{2}}y_{e}(\eta,\tau)}{|x-\eta|}\,|x-\eta|\,d\eta\right]^{2}dx$
$\hskip-85.35826pt\leq\,\frac{1}{h^{2}}\int_{x_{i-1}}^{x_{i}}\int_{x_{i-1}}^{x_{i}}|x-\eta|^{2}\,d\eta\int_{x_{i-1}}^{x_{i}}\frac{[\partial_{x_{2}x_{2}}y_{e}(x,\tau)-\partial_{x_{2}x_{2}}y_{e}(\eta,\tau)]^{2}}{|x-\eta|^{2}}\,d\eta dx$
$\hskip-85.35826pt\leq\,h\,\int_{x_{i-1}}^{x_{i}}\int_{x_{i-1}}^{x_{i}}\frac{[\partial_{x_{2}x_{2}}y_{e}(x,\tau)-\partial_{x_{2}x_{2}}y_{e}(\eta,\tau)]^{2}}{|x-\eta|^{2}}\,d\eta dx,$
we conclude
$\int_{x_{i-1}}^{x_{i}}[y(x,0)-I_{h_{2}}y(x,0)]^{2}dx$
$\hskip 28.45274pt\leq\,h^{4}\,\int_{x_{i-1}}^{x_{i}}\int_{x_{i-1}}^{x_{i}}\frac{[\partial_{x_{2}x_{2}}y_{e}(x,\tau)-\partial_{x_{2}x_{2}}y_{e}(\eta,\tau)]^{2}}{|x-\eta|^{2}}\,d\eta dx+h^{4}\int_{0}^{h}[\hat{y}_{i}(\tau)]^{2}d\tau$
$\hskip 28.45274pt\leq\,h^{4}\,|y_{e}|^{2}_{H^{5/2}((x_{i-1},x_{i})\times(0,h))}.$
When summing up all contrubutions, this finally gives
$\|y-P_{h}y\|_{L^{2}(\Gamma_{0})}\leq c\,h^{2}\,|y_{e}|_{H^{5/2}(\Omega)}.$
For all other boundary edges, the proof follows the same lines. ∎

### v2

@@
-Numerical solution of elliptic distributed optimal control problems with boundary value tracking
+
 Abstract
@@
 The boundary tracking is a special case of the partial tracking of a given target in a subset of $\Omega$ that is also called limited observation; see, e.g., [5].
-In this contribution, we consider the energy regularization corresponding to the choice $U=\widetilde{H}^{-1}(\Omega):=[H^{1}(\Omega)]^{*}$, where we use the standard notations for Lebesgue and Sobolev spaces; see, e.g., [8].
+There are many important practical applications, e.g., inverse heat transfer, photoacoustic tomography, etc., see also [4].
+In this contribution, we consider the energy regularization corresponding to the choice $U=\widetilde{H}^{-1}(\Omega):=[H^{1}(\Omega)]^{*}$ such that the solution of (2) defines an isomorphism when considering $y_{\varrho}\in H^{1}(\Omega)$.
+When using $U=L^{2}(\Omega)$ this implies $y_{\varrho}\in H^{1}_{\Delta}(\Omega):=\{y\in H^{1}(\Omega):\Delta y\in L^{2}(\Omega)\}$ to ensure an isomorphism, which afterwards requires higher order basis functions for a conforming discretization.
+Here we use the standard notations for Lebesgue and Sobolev spaces; see, e.g., [8].
 In order to follow the abstract theory presented in [3], we define the state space in such a way that the state-to-control map is an isomorphism.
@@
 Here we restrict the analysis to a conforming tensor-product finite element (fe) discretization that finally leads to a linear system of algebraic equations for which fast solvers can be constructed.
+Note that $c$ denotes a universal constant independent of the regularization parameter $\varrho$, and the discretization parameter $h$.
 2 State-based variational reformulation
-The variational formulation of the Neumann bvp (2) reads to find $y_{\varrho}\in H^{1}(\Omega)$ such that
+The variational formulation of the Neumann boundary value problem (2) reads to find $y_{\varrho}\in H^{1}(\Omega)$ such that
 $\langle\nabla y_{\varrho},\nabla y\rangle_{L^{2}(\Omega)}+\langle y_{\varrho},y\rangle_{L^{2}(\Omega)}=\langle u_{\varrho},y\rangle_{\Omega}$ (3)
 is satisfied for all $y\in H^{1}(\Omega)$, where we assume $u_{\varrho}\in\widetilde{H}^{-1}(\Omega)$.
+The variational problem has a unique solution due to Lax-Milgram’s lemma; see, e.g., [8].
 While the Neumann boundary condition in (2) enters the variational formulation (3) in a natural way, this condition has to be included in the definition of the state space
@@
 $\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}\leq c\,\varrho\,\|\overline{y}\|_{H^{1}(\Gamma)},\quad\|y_{\varrho}-\overline{y}_{e}\|_{H^{1}(\Omega)}\leq c\,\varrho^{1/2}\,\|\overline{y}\|_{H^{1}(\Gamma)}.$ (8)
-Note that the extension $\overline{y}_{e}\in Y$ of $\overline{y}\in H^{1/2}(\Gamma)$ as used in Lemma 2.1 can be computed as solution of the first boundary value problem for the biharmonic equation,
-$\Delta^{2}\overline{y}_{e}=0\quad\mbox{in}\;\Omega,\quad\overline{y}_{e}=\overline{y},\quad\partial_{n}\overline{y}_{e}=0\quad\mbox{on}\;\Gamma.$ (9)
+Here, the constant $c$ depends on the mapping properties of the partial differential operator $I-\Delta$ in certain Sobolev spaces, and the trace theorem.
 3 Conforming FE discretization on tensor product meshes
@@
 Let $y_{\varrho h}\in Y_{h}$ be the unique solution of (9). Then, for $\overline{y}\in L^{2}(\Gamma)$, there holds the error estimate
-$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq(1+\sqrt{2})\,\|\overline{y}\|_{L^{2}(\Gamma)},$ (12)
-while, for $\overline{y}\in H^{1/2}(\Gamma)$ and choosing $\varrho=h$, we have
+$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq(1+\sqrt{2})\,\|\overline{y}\|_{L^{2}(\Gamma)}.$ (11)
+If $\overline{y}\in H^{1/2}(\Gamma)$ is the Dirichlet trace of $\overline{y}_{e}\in Y$, and choosing $\varrho=h$, we have
 $\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq c\,h^{1/2}\,\|\overline{y}\|_{H^{1/2}(\Gamma)}.$ (12)
-A related result holds for $\overline{y}\in H^{1}(\Gamma)$ where, for $\varrho=h$,
+For $\overline{y}=\overline{y}_{e|\Gamma}\in H^{1}(\Gamma)$, $\overline{y}_{e}\in H^{3/2+\varepsilon}(\Omega)$ with $\varepsilon>0$, and choosing $\varrho=h$, we have
 $\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq c\,h\,\|\overline{y}\|_{H^{1}(\Gamma)}.$ (13)
@@
 $\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}\leq\|y_{\varrho h}-y_{\varrho}\|_{L^{2}(\Gamma)}+\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}\leq(1+\sqrt{2})\,\|\overline{y}\|_{L^{2}(\Gamma)}\,.$
-Next we consider the case when $\overline{y}_{e}\in Y$ is the extension of $\overline{y}$ as solution of the biharmonic boundary value problem (9).
-When using the triangle inequality twice in combination with Cea’s estimate (11), we can write, for arbitrary $y_{h}\in Y_{h}$
+For $\overline{y}=\overline{y}_{e|\Gamma}\in H^{1/2}(\Gamma)$ we use the triangle inequality twice and Cea’s estimate (10) for arbitrary $y_{h}\in Y_{h}$ to write
 $\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}^{2}$ $\leq$ $2\,\|y_{\varrho}-\overline{y}\|_{L^{2}(\Gamma)}^{2}+2\,\|y_{\varrho}-y_{\varrho h}\|^{2}_{L^{2}(\Gamma)}$
@@
 + $\hskip 56.9055pt+\,4\,\varrho\,\|y_{\varrho}-\overline{y}_{e}\|_{H^{1}(\Omega)}^{2}+4\,\varrho\,\|\overline{y}_{e}-y_{h}\|^{2}_{H^{1}(\Omega)}.$
-In particular for $y_{h}=P_{h}\overline{y}_{e}\in Y_{h}$ being the $L^{2}$ projection of $\overline{y}_{e}$, we have the standard fe error estimates
+In particular, for $y_{h}=P_{h}\overline{y}_{e}\in Y_{h}$ being the $L^{2}$ projection of $\overline{y}_{e}$, we have the standard fe error estimates; see, e.g., [8],
 $\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{H^{1}(\Omega)}\leq c\,\|\overline{y}_{e}\|_{H^{1}(\Omega)},\quad\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{L^{2}(\Omega)}\leq c\,h\,\|\overline{y}_{e}\|_{H^{1}(\Omega)}.$ (15)
@@
 and (12) follows when choosing $\varrho=h$.
-Finally, we consider the case $\overline{y}\in H^{1}(\Gamma)$ which implies $\overline{y}_{e}\in H^{3/2}(\Omega)$. In this case, we have
+Finally, we consider the case $\overline{y}=\overline{y}_{e|\Gamma}\in H^{1}(\Gamma)$ with $\overline{y}_{e}\in H^{3/2+\varepsilon}(\Omega)$ for some $\varepsilon>0$. In this case, we have
 $\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{H^{1}(\Omega)}\leq c\,h^{1/2}\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)},\;\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{L^{2}(\Omega)}\leq c\,h^{3/2}\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)},$
-and
-$\|\overline{y}-P_{h}\overline{y}_{e}\|_{L^{2}(\Gamma)}\leq c\,h\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)}.$
-Together with (8) we therefore conclude
-$\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}^{2}\,\leq\,c\,(\varrho^{2}+\varrho h+h^{2})\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)}^{2},$
-and for $\varrho=h$ this gives (14). ∎
+and   $\|\overline{y}-P_{h}\overline{y}_{e}\|_{L^{2}(\Gamma)}\leq c\,h\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)}.$ Together with (8) we therefore conclude $\|y_{\varrho h}-\overline{y}\|_{L^{2}(\Gamma)}^{2}\,\leq\,c\,(\varrho^{2}+\varrho h+h^{2})\,\|\overline{y}_{e}\|_{H^{3/2}(\Omega)}^{2},$ and, for $\varrho=h$, this gives (13).
+∎
 While the regularization error estimates as given in Lemma 2.1 are optimal in $\varrho$ for $\overline{y}\in H^{1}(\Gamma)$, i.e., $\overline{y}_{e}\in H^{3/2}(\Omega)$, we can expect higher order convergence for the fe approximation $P_{h}\overline{y}_{e}$ when $\overline{y}_{e}$ is more regular.
@@
 $\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{H^{1}(\Omega)}\leq c\,h\,\|\overline{y}_{e}\|_{H^{2}(\Omega)},\;\|\overline{y}_{e}-P_{h}\overline{y}_{e}\|_{L^{2}(\Omega)}\leq c\,h^{2}\,\|\overline{y}_{e}\|_{H^{2}(\Omega)}.$
-Since $\overline{y}_{e}$ is defined as solution of the biharmonic boundary value problem (9), we have to formulate additional compatibility conditions in order to ensure $\overline{y}_{e}\in H^{2}(\Omega)$, when $\Gamma=\partial\Omega$ is piecewise smooth.
+Note that $\overline{y}_{e}\in H^{2}(\Omega)$ implies some additional compatibility conditions on $\overline{y}$, when $\Gamma=\partial\Omega$ is piecewise smooth.
 In particular, the tangential derivatives of $\overline{y}$ have to be in $H^{1/2}(\Gamma)$, see, e.g., the discussion in [2, 6] in the two-dimensional case.
@@
 The matrices ${\mathbf{M}}_{BB}$, $\widetilde{\mathbf{K}}_{BB}$ and $\mathring{\mathbf{K}}_{h}$ are defined by the identities $({\mathbf{M}}_{BB}\mathbf{y}_{B},\mathbf{v}_{B})=\langle y_{h},v_{h}\rangle_{L^{2}(\Gamma)}$, $(\widetilde{\mathbf{K}}_{BB}\mathbf{y}_{B},\mathbf{v}_{B})=\langle y_{h},v_{h}\rangle_{H^{1}(\widetilde{\Omega}_{h})}$, $(\mathring{\mathbf{K}}_{h}\mathbf{y}_{h},\mathbf{v}_{h})=\langle y_{h},v_{h}\rangle_{H^{1}(\mathring{\Omega}_{h})}$ for all $\mathbf{y}_{h}=(\mathbf{y}_{I}^{\top},\mathbf{y}_{B}^{\top})^{\top}\leftrightarrow y_{h},v_{h}\in Y_{h}$ (fe isomorphism), where $\mathring{\Omega}_{h}=\Omega\setminus\overline{\widetilde{\Omega}}_{h}=(h,1-h)^{d}$ and $\widetilde{\Omega}_{h}=\Omega\setminus\overline{\mathring{\Omega}}_{h}$, whereas $\mathbf{\overline{y}}_{h}=(\mathbf{0}_{I}^{\top},\mathbf{\overline{y}}_{B}^{\top})^{\top}\in\mathbb{R}^{m}$ is given by $(\mathbf{\overline{y}}_{B},\mathbf{y}_{B})=\langle\overline{y}_{h},y_{h}\rangle_{L^{2}(\Gamma)}$ for all $\mathbf{y}_{h}=(\mathbf{y}_{I}^{\top},\mathbf{y}_{B}^{\top})^{\top}\leftrightarrow y_{h},v_{h}\in Y_{h}$.
-Eliminating $\mathbf{y}_{I}=-\mathring{\mathbf{K}}_{II}^{-1}\mathring{\mathbf{K}}_{IB}\mathbf{y}_{B}$ from the linear system (19), we arrive at the boundary Schur complement system
+Eliminating $\mathbf{y}_{I}=-\mathring{\mathbf{K}}_{II}^{-1}\mathring{\mathbf{K}}_{IB}\mathbf{y}_{B}$ from the linear system (18), we arrive at the boundary Schur complement (SC) system
 ${\mathbf{S}}_{BB}\mathbf{y}_{B}=\mathbf{\overline{y}}_{B}$ (19)
@@
 ${\mathbf{M}}_{BB}\leq{\mathbf{S}}_{BB}={\mathbf{M}}_{BB}+\varrho(\widetilde{\mathbf{K}}_{BB}+\mathring{\mathbf{S}}_{BB})\leq(1+\widetilde{c}\varrho h^{-1}+\mathring{c}\varrho h^{-1}){\mathbf{M}}_{BB}$ (20)
-with $h$ and $\varrho$ independent positive constants $\widetilde{c}$ and $\mathring{c}$.
+with $h$ and $\varrho$ independent positive constants $\widetilde{c}$ and $\mathring{c}$ arising from the estimates $\lambda_{max}(\mathbf{M}_{BB}^{-1}\widetilde{\mathbf{K}}_{BB})\leq\widetilde{c}h^{-1}$ and $\lambda_{max}(\mathbf{M}_{BB}^{-1}\mathring{\mathbf{S}}_{BB})\leq\mathring{c}h^{-1}$ of the maximal eigenvalues of $\mathbf{M}_{BB}^{-1}\widetilde{\mathbf{K}}_{BB}$ and $\mathbf{M}_{BB}^{-1}\mathring{\mathbf{S}}_{BB}$, respectively.
 The choice $\varrho\leq h$ delivers the desired result.
@@
 Table 1 provides the numerical results starting from level $\ell=1$ with $27$ dofs and running to the finest discretization level $\ell=7$ obtained by $6$ uniform refinements of the initial mesh.
-The fourth column displays the $L^{2}$ error $\|y-\overline{y}\|_{L^{2}(\Gamma)}$ on the boundary.
+The fourth column displays the $L^{2}$ error $\|y_{\ell}-\overline{y}\|_{L^{2}(\Gamma)}$ between the computed fe solution $y_{\ell}=y_{\varrho_{\ell}h_{\ell}}$ and the target $\overline{y}$ on the boundary.
 As expected, we observe second order of convergence; cf.
@@
 77 $16,581,375$ $2^{-8}$ 4.6054.605e5-5 $2.02$ 44 $29$ 44
-Table 1: Target (22): error = $\|y-\overline{y}\|_{L^{2}(\Gamma)}$, number of AMG CG iterations (#ACG-PCG its) for the original system (19), and number of CG (#SCG its) and lumped-mass preconditioned CG (#PSCG its) iterations for the Schur complement system (20), $\varrho=h^{2}$.
+Table 1: Target (21): error = $\|y_{\ell}-\overline{y}\|_{L^{2}(\Gamma)}$, eoc = $\log_{2}(\|y_{\ell}-\overline{y}\|_{L^{2}(\Gamma)}/\|y_{\ell+1}-\overline{y}\|_{L^{2}(\Gamma)})$, number of AMG CG iterations (#AMG-PCG its) for the original system (18), and number of CG (#SCG its) and lumped-mass preconditioned CG (#PSCG its) iterations for the SC system (19), $\varrho=h^{2}$.
 The second target
@@
 is a trace of a smooth function which does not fulfill the homogeneous Neumann boundary conditions on the boundary $\Gamma$ of $\Omega=(0,1)^{3}$, i.e., $\overline{y}\in H^{3/2-\varepsilon}(\Gamma)$, and $\overline{y}_{e}\in H^{2-\varepsilon}(\Omega)$, $\varepsilon>0$.
-Similar as in (17) we therefore expect a reduced eco of about 1.5.
+Similar as in (16) we therefore expect a reduced eoc of about 1.5.
 We perform the same tests for the target (22) as for the previous example for the target (21).
@@
 Table 2: Target (22): Same agenda as Table 1, but now $\varrho=h^{3/2}$.
+Finally, we consider a discontinuous target $\overline{y}$ that is equal to 1 inside the region $0.25,0.75]\times[0.25,0.75]$ and 0 outside on each face $\Gamma_{i}\subset\Gamma=\partial\Omega,i=1,\dots,6$ the unit cube $\Omega=(0,1)^{3}$.
+We perform the same tests as for the previous examples.
+The results are given in Table 3, where we have used $\varrho=h$ as prescribed by the theory.
+We observe the expected convergence rate eoc = 0.50.
+Furthermore, the iteration numbers are again independent of $h$, and show the same behavior as in the case of the preceding two examples.
+Figure 2 shows the computed state $y_{\ell}$ on the surface $\Gamma$ of the cube $\Omega$ (left) and on three cuts through the cube.
+$\ell$ #Dofs $h$ error eoc #AMG-PCG its #SCG its #PSCG its
+11 $27$ $2^{-2}$ 9.8179.817e2-2 - 33 33 33
+22 $343$ $2^{-3}$ 6.5356.535e1-1 - 44 $10$ $10$
+33 $3,375$ $2^{-4}$ 4.7684.768e1-1 $0.45$ 55 $17$ $17$
+44 $29,791$ $2^{-5}$ 3.5303.530e1-1 $0.43$ 44 $22$ 1919
+55 $250,047$ $2^{-6}$ 2.5402.540e1-1 $0.47$ 44 $23$ 2121
+66 $2,048,383$ $2^{-7}$ 1.8091.809e1-1 $0.49$ 44 $23$ 2121
+77 $16,581,375$ $2^{-8}$ 1.2831.283e1-1 $0.50$ 55 $23$ 2121
+Table 3: Discontinuous target with 1 inside the region of $0.25\quad 0.75]\times[0.25\quad 0.75]$ and 0 outside this region on each face of the unique cube: Same agenda as Table 1, but now $\varrho=h$.
+Figure 2: Computed state for the discontinuous target at level 77.
 6 Conclusions and Outlook
@@
 Lemma 6.4.
-Let $\Gamma=\partial\Omega$ for $\Omega=(0,1)^{d}$.
-Assume $y\in H^{2}(\Gamma)$, i.e., there exists an extension $y_{e}\in Y\cap H^{5/2}(\Omega)$ as solution of the biharmonic boundary value problem (9).
-Then, there holds the error estimate
+Let $\Gamma=\partial\Omega$ for $\Omega=(0,1)^{d}$. Assume $y\in H^{2}(\Gamma)$, i.e., there exists an extension $y_{e}\in Y\cap H^{5/2}(\Omega)$. Then, there holds the error estimate
 $\|y-P_{h}y\|_{L^{2}(\Gamma)}\leq c\,h^{2}\,|y|_{H^{2}(\Gamma)}.$ (34)